# Graph Learning

### v1

@@ Abstract
Graph learning has emerged as a pivotal artificial intelligence (AI) solution, driven by the increasing recognition of graph structures in modeling complex relationships within data [xia2021graph_GLsurvey, DBLP:journals/tbd/ZhangYZZ20, shehzad2024graph]. Graphs, which consist of nodes and edges, are inherently suited to represent a diverse range of complex systems such as social systems [zhou2023hierarchical_socialNetworks, yucage], knowledge graphs [peng2023knowledge_knowledgeGraph, 10880109], academic networks [liu2019shifu2, liu2022deep], urban dynamics [kong2022exploring_transportation, 10992276], and biological networks [yu2024long, peng2025biologically]. Unlike other Euclidean data such as image and text, graphs are typically non-Euclidean data, which exhibit irregular structures, such as varying node degrees and edge weights (see Fig. 1.1).
Graph learning techniques are specifically designed to leverage the structural information embedded in graphs. Besides, graph learning is not limited to provide solutions for graph-structured data, but also Euclidean data [DBLP:conf/cvpr/BrynteIOK24]. By extracting nodes and relations from Euclidean data, the graph structure can be first constructed and then be represented by graph learning solutions. For example, image segmentation and object detection can be implemented with graph learning by constructing a connection graph between pixels [DBLP:conf/aaai/HeWZZL22, DBLP:conf/cvpr/BrynteIOK24]. Moreover, graph learning can be applied to disease prediction by constructing a similarity graph of symptoms between patients. The efficiency of analyzing graphs is closely tied to how they are represented.
@@
Generally speaking, graph learning refers to machine learning on graphs, and its target is to extract the desired feature of a graph. Most graph learning approaches utilize deep learning techniques to encode and represent graph data as vectors in a continuous space. This representation allows graphs to be seamlessly integrated into downstream tasks, making graph learning a highly effective tool for various AI applications.
As illustrated in Figure 1.2, existing graph learning methods fall roughly into the following four categories [xia2021graph_GLsurvey]: methods based on deep learning, matrix factorization, random walk, and graph signal processing (GSP). Deep learning based methods include, for example, graph convolutional networks (GCNs), graph attention networks (GATs), graph auto-encoders (GAEs), graph generative networks, and graph spatial-temporal networks  [DBLP:journals/tcyb/PanHFLJZ20, DBLP:journals/corr/BrunaZSL13, DBLP:conf/nips/VaswaniSPUJGKP17, luan2024heterophilic]. Matrix factorization techniques can be categorized into two main types: graph Laplacian matrix factorization and vertex proximity matrix factorization  [choquette2024amplified, DBLP:books/ox/07/GolubR07, wan2023z, li2025revisiting]. Random walk-based methods include, for example, structure-based random walks, random walks that incorporate both structure and node information, random walks in heterogeneous networks, and those in time-varying networks  [DBLP:conf/ijcai/YangLZSC15, DBLP:journals/ml/LaoC10, nguyen2018continuous_CDNE, xia2019random_randomWalk]. Finally, GSP focuses on the sampling and recovery of graphs, as well as inferring the topological structure of data  [levie2024graphon, DBLP:journals/tsp/SandryhailaM13, DBLP:journals/tsp/ChenVSK15, DBLP:conf/icassp/NarangGO13]. It is worth noting that graph learning [xia2021graph_GLsurvey] and graph neural networks (GNNs) [corso2024graph, zhou2022graph] are not exactly the same thing, with GNNs being a popular type of graph learning approaches.
@@
To better illustrate diverse graph learning approaches, we start with a classic task: community detection. In social networks, community detection aims to identify groups of users who interact more frequently with each other than with those outside their group. We frame the community detection task using the following graph learning methods below.
@@
    $h_{v}^{new}=\sigma\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{u\in N(v)}\alpha_{uv}^{k}h_{u}^{\textquoteright}\right),$ (1.2)
    where $N(v)$ represents the neighbors of $v$, $K$ represents the number of heads, $\alpha_{uv}^{k}$ is the normalizing attention coefficients, and $h_{u}^{\textquoteright}$ is the original feature vector. This iterative process allows the model to focus on the most informative connections, thereby enhancing its performance.
    GAEs contain encoders and decoders. In GAEs, GNNs are employed as encoders to embed networks vertices into low-dimensional vectors, and decoders convert these vectors into reconstructed graph data. Specific auto-encoder for graphs are designed by combining either GCN with Generative Adversarial Network (GAN) or Long Short-Term Memory (LSTM) with GAN. GAEs can learn both, topology structure and content features  [DBLP:conf/aaai/CaoLX16, DBLP:conf/kdd/WangC016]. The loss function of GAEs is the mean squared error loss represented as:
@@
  • Matrix factorization based methods: Matrix factorization techniques utilize spectral graph theory to examine community structures within graphs [wan2023z]. A key tool in this analysis is the Laplacian matrix $L$ , defined as $L=D-A$ , where $D$ represents the degree matrix and $A$ is the adjacency matrix. The eigenvalues and eigenvectors of $L$ offer valuable insights into the graph’s connectivity. In spectral clustering, the eigenvectors associated with the smallest eigenvalues are particularly informative, as they can reveal the underlying community structure.
@@
  • Random walks based methods: Random walks play a crucial role in understanding the flow of information across the graph [xia2019random_randomWalk]. Random walk-based methods classified as structure and node information-based random walk methods utilize random walks to generate node embeddings to capture local and global structures. Algorithms such as DeepWalk [perozzi2014deepwalk] and Node2Vec [grover2016node2vec] work in alignment with such techniques. By optimizing a skip-gram model over these walks, nodes that frequently appear together in walks are embedded closely in the vector space, allowing for effective clustering based on proximity in the embedding space. These structure-based random-walk methods leverage structural information in networks for community detection.
    In addition to network topology, content or label of vertex also improve the accuracy of representation, eventually guiding the random walk process. Structure and node information-based random walks methods regard vertex attribute information as supplementary information. For example, TADW [DBLP:conf/ijcai/YangLZSC15] and MMDW [tu2016max] are algorithms based on DeepWalk, leveraging text and labeling information of vertices to enhance the performance, respectively. Thus, the transition probabilities of moving from one node to another are adjusted based on both the structural properties of networks and the attributes of vertex.
@@
Graph learning has advanced significantly in recent years, yet several critical challenges remain. For instance, scalability remains a primary concern, as real-world graphs, like social networks or biological systems, often contain billions of nodes and edges, straining computational resources and necessitating efficient algorithms for processing large-scale data [ju2024survey]. Dynamic and temporal graphs [febrinanto2023graph_GLL], which evolve over time, introduce complexity in modeling time-dependent relationships and maintaining real-time updates, critical for applications like financial fraud detection. Multimodal graphs, integrating diverse data types (e.g., text, images, and numerical features), require robust methods to align and fuse heterogeneous information effectively [peng2024learning]. In parallel, the growing influence of generative AI opens new opportunities and risks for graph learning [li2025generative], particularly in generating realistic graphs or augmenting sparse data, but ensuring validity and fidelity remains difficult. Explainability is another pressing issue [huang2025sehg, nandan2025graphxai], as the black-box nature of models like GNNs obscures decision-making, limiting trust in high-stakes domains like healthcare. This ties closely with responsible AI concerns, including, e.g., fairness, robustness, and privacy. For instance, bias in graph structures or node attributes can propagate through models [luo2024algorithmic], while sensitive relational data demands privacy-preserving graph learning techniques. Addressing these challenges is essential for deploying graph learning systems that are not only effective but also trustworthy and socially responsible.
@@ 2.1 Graph Data Summarization
This section categorizes the graph data summarization techniques into three major groups: graph sparsification, condensation, and coarsening. Figure 2.3 compares the three graph data summarization strategies. Graph sparsification selects a subset of nodes and edges from the original graph to preserve critical properties, such as essential edges or node centrality [chen2023demystifying_sparsificationSimplest]. On the other hand, graph coarsening, groups and aggregates similar nodes into super-nodes and their connections into super-edges instead of removing nodes [chen2023gromov_GWGRAPH]. Graph condensation [sun2024gc_graphCondensationBenchmark], a more recent method, differs from sparsification and coarsening by creating a new, smaller synthetic graph that retains the task performance of graph learning while reducing graph size. We discuss each of the techniques in the sub-sections below:
@@ 2.1.2 Graph Coarsening
Graph coarsening is a promising direction for improving the scalability of graph processing by summarizing a very large graph into a smaller, tractable one. The selection of nodes or edges in sparsification methods can lead to the loss of information. Graph coarsening ensures that sufficient valuable information from the graph structure is preserved by grouping and aggregating nodes or edges. Given $G=(V,E)$ graph coarsening produce new graph $G^{\prime}=(V^{\prime},E^{\prime})$ which contains super-nodes $V^{\prime}$ and super-edges $E^{\prime}$, where $|V^{\prime}|<|V|$ and $|E^{\prime}|<|E|$. The coarse graph should be tractable and mapped from the original graph, which can be represented by a one-hot matrix $C\in\{0,1\}^{|V|\times|V^{\prime}|}$ to denote the membership of a group of nodes. The adjacency matrix of the coarsened graph is $A^{\prime}=C^{\intercal}AC$, and the Laplacian matrix is $L^{\prime}=C^{\intercal}LC$.
@@
Classical Coarsening Techniques. The algebraic multigrid literature has influenced the development of graph coarsening [chen2023gromov_GWGRAPH]. The classical graph coarsening techniques include some techniques that use traditional graph partitioning, clustering, and spectral properties to group nodes into supernodes while reducing graph size. One example of clustering is spectral clustering [jin2020graph_spectralClustering], which leverages the eigenvectors of the graph Laplacian to find dense partitions within the graph, summarizing its structure. An earlier technique, Edge Matching, introduced by dhillon2007weighted_edgeMatching, merges node pairs by maximizing their total edge weight. This approach typically computes the weight for each pair of nodes $(v_{i},v_{j})$, defined as $\frac{W_{ij}}{\max(d_{i},d_{j})}$, where $d_{i}$ and $d_{j}$ represent the degrees of the nodes. This ensures that significant connections in the graph are preserved during the reduction process. Kron Reduction [dorfler2012kron] is an algebraic method that reduces a graph by applying the Schur complement to its Laplacian matrix, preserving essential physical properties like effective resistance, which is critical in applications such as electrical networks. CoarseNet [purohit2014fast_coarseNet] focuses on retaining diffusion properties by grouping nodes based on their influence within the network, enabling substantial graph reduction without losing key information on how processes propagate. Compressive network embedding (COSINE) by zhang2020cosine_cosine performs coarsening by partitioning nodes into groups for parameter sharing in compressive embeddings. deng2020graphzoom introduced GraphZoom uses spectral properties to aggregate nodes with high similarity and builds multi-level coarsened graphs to improve the scalability of graph embedding tasks. fahrbach2020faster_schurComplement used Schur Complement and Random Contraction techniques to eliminate or combine nodes into supernodes, maintaining the graph’s spectral and structural properties to speed up embedding and simplify graph structure. Similarly, featured graph coarsening (FGC) by kumar2023featured_FGC compresses both node features and graph structure to preserve the node relationships, which is particularly useful for tasks like node classification.
@@ 2.1.3 Graph Condensation
$\mathcal{S}=\min_{\mathcal{S}}\mathcal{L}^{\mathcal{T}}\left(\theta^{\mathcal{S}}\right)\quad\text{s.t.}\quad\theta^{\mathcal{S}}=\arg\min_{\theta}\mathcal{L}^{\mathcal{S}}(\theta),$ (2.3)
@@
$\ell_{\text{MMD}}=\sum_{c\in C}r_{c}\cdot\left|\text{Mean}(E_{c})-\text{Mean}(\tilde{E}_{c})\right|^{2}\end{aligned}.$ (2.5)
@@ 2.2.1 Node-wise Sampling
In traditional graph learning methods, node-wise sampling is often applied in algorithms like random walks [xia2019random_randomWalk], PageRank [brin1998pagerank], DeepWalk [perozzi2014deepwalk], and Node2Vec [grover2016node2vec], which focus on selecting individual nodes and their direct neighborhoods in the process of graph traversal or exploration and to learn graph representations. In contrast, in GNNs, node-wise sampling specifically aims to reduce the computational load during the aggregation of neighbor features.Random Walk traverses a graph by moving from one node to a randomly selected neighbor, capturing both local and global connectivity. PageRank builds on this concept by introducing a transition mechanism that allows the walker to jump to random nodes, assigning importance to nodes based on their visit probabilities and sampling them accordingly. DeepWalk refines this approach by performing truncated random walks, limiting the number of steps to balance computational efficiency while generating node sequences. These sequences are then treated like sentences in a language model and embedded into a vector space using the Skip-Gram model. Node2Vec further enhances DeepWalk by incorporating a biased random walk strategy that blends breadth-first search (BFS) and depth-first search (DFS). This hybrid approach enables the model to capture both local neighborhood structures and long-range dependencies, leading to more informative node representations.

### v2

@@ Abstract
Graph learning has emerged as a pivotal artificial intelligence (AI) solution, driven by the increasing recognition of graph structures in modeling complex relationships within data [xia2021graph_GLsurvey, DBLP:journals/tbd/ZhangYZZ20, shehzad2024graph]. Graphs, which consist of nodes and edges, are inherently suited to represent a diverse range of complex systems such as social systems [zhou2023hierarchical_socialNetworks, yucage], knowledge graphs [peng2023knowledge_knowledgeGraph, 10880109], academic networks [liu2019shifu2, liu2022deep], urban dynamics [kong2022exploring_transportation, 10992276], and biological networks [yu2024long, peng2025biologically]. Unlike other Euclidean data such as images and text, graphs are typically non-Euclidean data, which exhibit irregular structures, such as varying node degrees and edge weights (see Fig. 1.1).
Graph learning techniques are specifically designed to leverage the structural information embedded in graphs. Besides, graph learning is not limited to providing solutions for graph-structured data, but also Euclidean data [DBLP:conf/cvpr/BrynteIOK24]. With nodes and relations extracted from Euclidean data, the graph structure can be first constructed and then be represented by graph learning solutions. For example, image segmentation and object detection can be implemented with graph learning by constructing a connection graph between pixels [DBLP:conf/aaai/HeWZZL22, DBLP:conf/cvpr/BrynteIOK24]. Moreover, graph learning can be applied to disease prediction by constructing a similarity graph of symptoms between patients. The efficiency of analyzing graphs is closely tied to how they are represented.
@@
Generally speaking, graph learning refers to machine learning on graphs, and its target is to extract the desired features of a graph. Most graph learning approaches utilize deep learning techniques to encode and represent graph data as vectors in a continuous space. This representation allows graphs to be seamlessly integrated into downstream tasks, making graph learning a highly effective tool for various AI applications.
As illustrated in Figure 1.2, existing graph learning methods fall roughly into the following four categories [xia2021graph_GLsurvey]: methods based on deep learning, matrix factorization, random walk, and graph signal processing (GSP). Deep learning-based methods include, for example, graph convolutional networks (GCNs), graph attention networks (GATs), graph auto-encoders (GAEs), graph generative networks, and graph spatial-temporal networks [DBLP:journals/tcyb/PanHFLJZ20, DBLP:journals/corr/BrunaZSL13, DBLP:conf/nips/VaswaniSPUJGKP17, luan2024heterophilic]. Matrix factorization techniques can be categorized into two main types: graph Laplacian matrix factorization and vertex proximity matrix factorization [choquette2024amplified, DBLP:books/ox/07/GolubR07, wan2023z, li2025revisiting]. Random walk-based methods include, for example, structure-based random walks, random walks that incorporate both structure and node information, random walks in heterogeneous networks, and those in time-varying networks [DBLP:conf/ijcai/YangLZSC15, DBLP:journals/ml/LaoC10, nguyen2018continuous_CDNE, xia2019random_randomWalk]. Finally, GSP focuses on the sampling and recovery of graphs, as well as inferring the topological structure of data [levie2024graphon, DBLP:journals/tsp/SandryhailaM13, DBLP:journals/tsp/ChenVSK15, DBLP:conf/icassp/NarangGO13]. It is worth noting that graph learning [xia2021graph_GLsurvey] and graph neural networks (GNNs) [corso2024graph, zhou2022graph] are not exactly the same thing, with GNNs being a popular type of graph learning approaches.
@@
To better illustrate diverse graph learning approaches, we start with a classic task: community detection. In social networks, community detection aims to identify groups of users who interact more frequently with each other than with those outside their group. We frame the community detection task using the following graph learning methods.
@@
    $h_{v}^{new}=\sigma\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{u\in N(v)}\alpha_{uv}^{k}h_{u}^{’}\right),$ (1.2)
    where $N(v)$ represents the neighbors of $v$, $K$ represents the number of heads, $\alpha_{uv}^{k}$ is the normalized attention coefficients, and $h_{u}^{’}$ is the original feature vector. This iterative process allows the model to focus on the most informative connections, thereby enhancing its performance.
    GAEs contain encoders and decoders. In GAEs, GNNs are employed as encoders to embed network vertices into low-dimensional vectors, and decoders convert these vectors into reconstructed graph data. Specific auto-encoders for graphs are designed by combining either GCN with Generative Adversarial Network (GAN) or Long Short-Term Memory (LSTM) with GAN. GAEs can learn both topology structure and content features [DBLP:conf/aaai/CaoLX16, DBLP:conf/kdd/WangC016]. The loss function of GAEs is the mean squared error loss represented as:
@@
  • Matrix factorization based methods: Matrix factorization techniques utilize spectral graph theory to examine community structures within graphs [wan2023z]. A key tool in this analysis is the Laplacian matrix $L$ , defined as L=DAL=D-A , where $D$ represents the degree matrix and $A$ is the adjacency matrix. The eigenvalues and eigenvectors of $L$ offer valuable insights into the graph’s connectivity. In spectral clustering, the eigenvectors associated with the smallest eigenvalues are particularly informative, as they can reveal the underlying community structure.
@@
  • Random walks based methods: Random walks play a crucial role in understanding the flow of information across the graph [xia2019random_randomWalk]. Random walk-based methods, which are classified as structure- and node information-based random walk methods, utilize random walks to generate node embeddings to capture local and global structures. Algorithms such as DeepWalk [perozzi2014deepwalk] and Node2Vec [grover2016node2vec] work in alignment with such techniques. By optimizing a skip-gram model over these walks, nodes that frequently appear together in walks are embedded closely in the vector space, allowing for effective clustering based on proximity in the embedding space. These structure-based random-walk methods leverage structural information in networks for community detection.
    In addition to network topology, content or label of vertex also improves the accuracy of representation, eventually guiding the random walk process. Structure and node information-based random walks methods regard vertex attribute information as supplementary information. For example, TADW [DBLP:conf/ijcai/YangLZSC15] and MMDW [tu2016max] are algorithms based on DeepWalk, leveraging text and labeling information of vertices to enhance the performance, respectively. Thus, the transition probabilities of moving from one node to another are adjusted based on both the structural properties of networks and the attributes of vertex.
@@
Graph learning has advanced significantly in recent years, yet several critical challenges remain. For instance, scalability remains a primary concern, as real-world graphs, like social networks or biological systems, often contain billions of nodes and edges, straining computational resources and necessitating efficient algorithms for processing large-scale data [ju2024survey]. Dynamic and temporal graphs [febrinanto2023graph_GLL], which evolve over time, introduce complexity in modeling time-dependent relationships and maintaining real-time updates, critical for applications like financial fraud detection. Multimodal graphs, integrating diverse data types (e.g., text, images, and numerical features), require robust methods to align and fuse heterogeneous information effectively [peng2024learning]. In parallel, the growing influence of generative AI opens new opportunities and risks for graph learning [li2025generative], particularly in generating realistic graphs or augmenting sparse data, but ensuring validity and fidelity remains difficult. Explainability is another pressing issue [huang2025sehg, nandan2025graphxai], as the black-box nature of models like GNNs obscures decision-making, limiting trust in high-stakes domains like healthcare. This ties closely with responsible AI concerns, including, e.g., fairness, robustness, and privacy. For instance, bias in graph structures or node attributes can propagate through models [luo2024algorithmic], while sensitive relational data demands privacy-preserving graph learning techniques. Addressing these challenges is essential for deploying graph learning systems that are not only effective but also trustworthy and socially responsible. Furthermore, we offer a summary table as shown in Table 1.1. It summarizes each branch, application domains, and representative methods.
@@
Table 1.1: Overview of All Sections
Taxonomy Application Domains Representative methods
Scalable
Graph Learning
Social Network Analysis,
Recommendation Systems
Graph Data Summarization
Computational Sampling Methods
Distributed Graph Learning
Temporal
Graph Learning
Real-time Traffic Forecasting,
Epidemic Spread Prediction
Spatiotemporal Graph Learning
Dynamic Graph Learning
Multimodal
Graph Learning
Multimedia Content Analysis Graph-driven Multimodal Learning
Learning on Multimodal Graphs
Generative
Graph Learning
Drug Discovery,
Molecular Material Design
Unconditional Generative Graph Learning
Conditional Generative Graph Learning
Explainable
Graph Learning
Medical Diagnostics,
Financial Risk Control
Post-hoc Explanation for Graph Learning
Self-explanatory Graph Learning
Responsible
Graph Learning
Credit Analysis,
Judicial Risk Assessment,
Public Policy Formulation
Privacy-Preserving Graph Learning
Fairness in Graph Learning
@@ 2.1 Graph Data Summarization
This section categorizes the graph data summarization techniques into three major groups: graph sparsification, condensation, and coarsening. Figure 2.3 compares the three graph data summarization strategies. Graph sparsification selects a subset of nodes and edges from the original graph to preserve critical properties, such as essential edges or node centrality [chen2023demystifying_sparsificationSimplest]. On the other hand, graph coarsening groups and aggregates similar nodes into super-nodes and their connections into super-edges instead of removing nodes [chen2023gromov_GWGRAPH]. Graph condensation [sun2024gc_graphCondensationBenchmark], a more recent method, differs from sparsification and coarsening by creating a new, smaller synthetic graph that retains the task performance of graph learning while reducing graph size. We discuss each of the techniques in the sub-sections below:
@@ 2.1.2 Graph Coarsening
Graph coarsening is a promising direction for improving the scalability of graph processing by summarizing a very large graph into a smaller, tractable one. The selection of nodes or edges in sparsification methods can lead to the loss of information. Graph coarsening ensures that sufficient valuable information from the graph structure is preserved by grouping and aggregating nodes or edges. Given $G=(V,E)$, graph coarsening produces new graph $G^{\prime}=(V^{\prime},E^{\prime})$ which contains super-nodes $V^{\prime}$ and super-edges $E^{\prime}$, where $|V^{\prime}|<|V|$ and $|E^{\prime}|<|E|$. The coarse graph should be tractable and mapped from the original graph, which can be represented by a one-hot matrix $C\in\{0,1\}^{|V|\times|V^{\prime}|}$ to denote the membership of a group of nodes. The adjacency matrix of the coarsened graph is $A^{\prime}=C^{\intercal}AC$, and the Laplacian matrix is $L^{\prime}=C^{\intercal}LC$.
@@
Classical Coarsening Techniques. The algebraic multigrid literature has influenced the development of graph coarsening [chen2023gromov_GWGRAPH]. The classical graph coarsening techniques include some techniques that use traditional graph partitioning, clustering, and spectral properties to group nodes into supernodes while reducing graph size. One example of clustering is spectral clustering [jin2020graph_spectralClustering], which leverages the eigenvectors of the graph Laplacian to find dense partitions within the graph, summarizing its structure. An earlier technique, Edge Matching, introduced by dhillon2007weighted_edgeMatching, merges node pairs by maximizing their total edge weight. This approach typically computes the weight for each pair of nodes $(v_{i},v_{j})$, defined as $\frac{W_{ij}}{\max(d_{i},d_{j})}$, where $d_{i}$ and $d_{j}$ represent the degrees of the nodes. This ensures that significant connections in the graph are preserved during the reduction process. Kron Reduction [dorfler2012kron] is an algebraic method that reduces a graph by applying the Schur complement to its Laplacian matrix, preserving essential physical properties like effective resistance, which is critical in applications such as electrical networks. CoarseNet [purohit2014fast_coarseNet] focuses on retaining diffusion properties by grouping nodes based on their influence within the network, enabling substantial graph reduction without losing key information on how processes propagate. Compressive network embedding (COSINE) by zhang2020cosine_cosine performs coarsening by partitioning nodes into groups for parameter sharing in compressive embeddings. deng2020graphzoom introduced GraphZoom, which uses spectral properties to aggregate nodes with high similarity and builds multi-level coarsened graphs to improve the scalability of graph embedding tasks. fahrbach2020faster_schurComplement used Schur Complement and Random Contraction techniques to eliminate or combine nodes into supernodes, maintaining the graph’s spectral and structural properties to speed up embedding and simplify graph structure. Similarly, featured graph coarsening (FGC) by kumar2023featured_FGC compresses both node features and graph structure to preserve the node relationships, which is particularly useful for tasks like node classification.
@@ 2.1.3 Graph Condensation
$\mathcal{S}=\min_{\mathcal{S}}\mathcal{L}^{\mathcal{T}}\left(\theta^{\mathcal{S}}\right)\quad\text{s.t.}\quad\theta^{\mathcal{S}}=\arg\min_{\theta}\mathcal{L}^{\mathcal{S}}(\theta).$ (2.3)
@@
$\begin{aligned} \ell_{\text{MMD}}=\sum_{c\in C}r_{c}\cdot\left|\text{Mean}(E_{c})-\text{Mean}(\tilde{E}_{c})\right|^{2}\end{aligned}.$ (2.5)
@@ 2.2.1 Node-wise Sampling
In traditional graph learning methods, node-wise sampling is often applied in algorithms like random walks [xia2019random_randomWalk], PageRank [brin1998pagerank], DeepWalk [perozzi2014deepwalk], and Node2Vec [grover2016node2vec], which focus on selecting individual nodes and their direct neighborhoods in the process of graph traversal or exploration and to learn graph representations. In contrast, in GNNs, node-wise sampling specifically aims to reduce the computational load during the aggregation of neighbor features.
Random Walk traverses a graph by moving from one node to a randomly selected neighbor, capturing both local and global connectivity. PageRank builds on this concept by introducing a transition mechanism that allows the walker to jump to random nodes, assigning importance to nodes based on their visit probabilities and sampling them accordingly. DeepWalk refines this approach by performing truncated random walks, limiting the number of steps to balance computational efficiency while generating node sequences. These sequences are then treated like sentences in a language model and embedded into a vector space using the Skip-Gram model. Node2Vec further enhances DeepWalk by incorporating a biased random walk strategy that blends breadth-first search (BFS) and depth-first search (DFS). This hybrid approach enables the model to capture both local neighborhood structures and long-range dependencies, leading to more informative node representations.